Session: Feb 2017 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Hartebeesthoek (Jonathan Quick at Thu Mar 09 11:06:53 2017) Minor failures: ANT SCHED
May have lost 5 minutes to a power failure, and target source set before the last scan on it started. No other know problems.

Badary (Lysakova, Nekrasov at Thu Mar 09 12:00:28 2017) Success:

Medicina (orfei at Thu Mar 09 14:07:31 2017) Success:

Yebes (Miguel Sanz at Thu Mar 09 14:27:43 2017) Minor failures: TSYS
No calibration until scan no0019

Zelenchukskaya (Kononenko V. at Thu Mar 09 14:53:55 2017) Success:

Svetloe (Slobojaninova N. at Thu Mar 09 14:54:10 2017) Success:

Effelsberg (Uwe at Fri Mar 10 07:50:25 2017) Success:
Quite some rain, but weather improved to the end. No known problems.

Noto (Salvo Buttaccio at Fri Mar 10 09:42:28 2017) Minor failures: RECO
Problems with mk5b recorder. Data recorded on flexbuff in vdif format

Onsala (Jun Yang at Tue Mar 14 15:34:24 2017) Success:
A cloudy day again.

Torun (Pawel Wolak at Fri Mar 17 11:14:23 2017) Success: